OpenStudy (anonymous):

I have two questions that is messing me up, please help Which of the following is the formula for the geometric sequence? 900, 300, 100, 33 1/3,... an = 900(3)n-1 an = 33 1/3 (3)n-1 an = 900(1/3)n-1 Which of the following is the formula for the geometric sequence? 8, -16, 32, -64,... an = 8(-2)n-1 an = 2(-8)n-1 an = -8(2)n-1

OpenStudy (agent0smith):

For the first one you need to find the common ratio, like we did in the last question. Then all you need to do is find the first term, which is easy (as it's given to you).

OpenStudy (anonymous):

divided 3(first question

OpenStudy (agent0smith):

So the common ratio is 1/3. What's the first term?

OpenStudy (anonymous):

term i think it would be the first number in the line so 900?

OpenStudy (agent0smith):

Correct. The formula for a geometric sequence is \[\Large a_n = a_1 r^{n-1}\] where r is the common ratio, a1 is the first term. So which formula is correct, with the numbers you found?

OpenStudy (anonymous):

the last one?

OpenStudy (agent0smith):

Yep, an = 900(1/3)n-1 looks right.

OpenStudy (agent0smith):

The second one can be done in the same way. Find r, and a1.

OpenStudy (anonymous):

okay just a sec.

OpenStudy (agent0smith):

Okay, I need to go, but I can help later if you still need it. But you should be fine, just find r as you did before, divide a term by the term before it, then find a1, and find the formula that matches them.

OpenStudy (anonymous):

I cant find the patteren

OpenStudy (agent0smith):

To find r, divide a term by the term before it: 8, -16, 32, -64

OpenStudy (agent0smith):

The first term is easy, just the first number in the pattern.

OpenStudy (anonymous):

@tcarroll010

OpenStudy (agent0smith):

-16/8 = r = ??

OpenStudy (anonymous):

-2

OpenStudy (agent0smith):

So that's r. Now find the first term.

OpenStudy (anonymous):

first term is 8

OpenStudy (anonymous):

the first one??

OpenStudy (agent0smith):

an = 8(-2)n-1 is correct.

OpenStudy (anonymous):

:D thank you so much I understand now so r is the patteren and an is the first term right?

OpenStudy (agent0smith):

Yep. Well a1 is the first term, an means the nth term. r is the common ratio or multiple from one term to the next.

OpenStudy (agent0smith):

You're welcome! :D

OpenStudy (anonymous):

\[\sqrt{6}(\sqrt{3}+5\sqrt{2})\]

OpenStudy (agent0smith):

Distribute first: \[\large \sqrt{6}(\sqrt{3}+5\sqrt{2}) =\sqrt{6}\sqrt{3}+5\sqrt{6}\sqrt{2}) = \] \[\large \sqrt{6*3}+5\sqrt{6*2} = \sqrt{18}+5\sqrt{12} =\] \[\large \sqrt{9*2}+5\sqrt{4*3} = \sqrt{9}\sqrt{2}+5\sqrt{4}\sqrt{3} = \] Try and simplify it now.